OpenStudy (anonymous):

Use the distributive property to simplify the expression below: 6x-10 divided by -2

OpenStudy (luigi0210):

Same as the last one: \[\frac{ 6x-10 }{ -2 }=\frac{ -2(5-3x) }{ -2 }\]